OpenStudy (anonymous):

what is the simplified form of x plus 1 over x squared plus x minus 6 divided by x squared plus 5x plus 4 over x minus 2 ?

OpenStudy (anonymous):

\[\frac{x+1}{x^2+x-6} \div \frac{x^2+5x+4}{x-2}\]factor trinomials; invert and multiply\[=\frac{x+1}{(x+3)(x-2)}*\frac{x-2}{(x+1)(x+4)}\]cancel common factors in numerator and denominator\[=\frac{1}{(x+3)(x+4)}\]polynomial form is the more simplified form (multiply--FOIL)\[=\frac{1}{x^2+7x+12}\]

OpenStudy (anonymous):

@mandolino what is the answer?

OpenStudy (yaya090600):

\[\frac{ 1 }{ (x+3) (x+4) }\]